OpenStudy (anonymous):

-x-2=1-2x HELP, STEPS ALSO

OpenStudy (anonymous):

-x+2x=1+2 x=3

OpenStudy (anonymous):

but what are the steps to get that?

OpenStudy (mertsj):

Add 2x to both sides first.

OpenStudy (anonymous):

Okay thanks

OpenStudy (mertsj):